Let p, (t) 6+t, P2(t) =t-3t, p3 (t) = 1 +t-2t. Complete parts (a) and (b) below. Use coordinate vectors to show that these polynomials form a basis for P2. What are the coordinate vectors corresponding to p, p2, and pa? P- Place these coordinate vectors into the columns fa matrix A. What can be said about the matrix A? O A. The matrix A forms a basis for R3 by the Invertible Matrix Theorem because all square matrices are row equivalent to I3. O B. The matrix A is invertible because it is row equivalent to I and therefore the row reduced columns of A forma basis for R3 by the Invertible Matrix Theorem. O C. The matrix A is invertible because it is row equivalent to I3 and therefore the null space of A, denoted Nul A, forms a basis for R3 by the Invertible Matrix Theorem. O D. The matrix A is invertible because it is row equivalent to l3 and therefore the original columns of A form a basis for R3 by the Invertible Matrix Theorem. How does this show that the polynomials form a basis for P2? O A. The polynom ials form a basis for P2 because of the isomorphism between R3 and P2. O B. The polynomials form a basis for P2 because the columns of A are linearly independent. O C. The polynomials form a basis for P2 because the matrix A is invertible. O D. The polynomials form a basis for P2 because any basis in Rn is also a basis in P Consider the basis B= P1, P2. P3 for P2. Find q in P2. given that [qlB 1 2
q(t)= (+ t+ (D2
Answer #1

a. We know that the set {1,t,t2} is the standard basis of P2 and dim(P2) = 3.

Let M=

6

0

1

1

0

0

0

1

1

0

1

0

1

-3

-2

0

0

1

It may be observed that the entries in the columns of A are the   scalar multiples of 1 and the coefficients of t,t2 in p1(t), p2(t) , p3(t) and 1, t,t2.

The RREF of M is

1

0

0

1/5

-3/5

-1/5

0

1

0

1/5

-13/5

-6/5

0

0

1

-1/5

18/5

6/5

This implies that the vectors/polynomials p1(t), p2(t) , p3(t) are linearly independent and the vectors/polynomials 1, t,t2 are linear combinations of p1(t), p2(t) , p3(t).

Hence the vectors/polynomials p1(t), p2(t) , p3(t) form a basis for P2.

The coordinate vectors corresponding to p1, p2 , p3 are (6,0,1)T, (0,1,-3)T and (1,1,-2)T respectively.

Option B is the correct answer.

Option B is the correct answer.

b. If [q]B = (-3,1,2)T, then q = -3(6+t2)+1(t-3t2)+2(1+t-2t2) = -16+3t-10t2.